# Positive traces on certain $\operatorname{SL}(2)$ Coulomb branches

### v1

@@ 2.3 Positive traces
$I(f_{k})=\int_{\varepsilon S^{1}}f_{k}(q^{k}z)z^{-\frac{km}{2}}\overline{f_{k}}(z^{-1})\omega(z)\frac{dz}{z}$
is positive for any $k$ such that $f_{k}$ is nonzero. Let $f=f_{k}$, this is a function with the poles of order at most one at ±$\pm q^{k},k\in\mathbb{Z}$. Then the only possible poles of $f(q^{k}z)\overline{f}(z^{-1})$ are poles of order at most two at ±$\pm q^{k},k\in\mathbb{Z}$. Since $\omega(1)=\omega(-1)=0$ and $\omega(z)=\omega(z^{-1})$, the function $\omega$ has a double root at 111 and 1-1- 1. Using the quasi-periodicity condition we get that $w$ has a double root at ±$\pm q^{k},k\in\mathbb{Z}$. It follows that $f(q^{k}z)\overline{f}(z^{-1})\omega(z)$ is a holomorphic function on $\mathbb{C}^{*}$. Hence we can write
$I(f)=\int_{\varepsilon S^{1}}f(q^{k}z)z^{-\frac{km}{2}}\overline{f}(z^{-1})\omega(z)\frac{dz}{z}=$ $\int_{q^{-\frac{k}{2}}S^{1}}f(q^{k}z)z^{-\frac{km}{2}}\overline{f}(z^{-1})\omega(z)\frac{dz}{z}=$
$q^{\frac{k^{2}m}{4}}$ $\int_{S^{1}}f(q^{\frac{k}{2}}t)\overline{f}(q^{\frac{k}{2}}t^{-1})t^{-\frac{km}{2}}\omega(q^{-\frac{k}{2}}t)\frac{dt}{t}.$
For $z\in S^{1}$ we have $f(q^{\frac{k}{2}}t)\overline{f}(q^{\frac{k}{2}}t^{-1})=\lvert f(q^{\frac{k}{2}}t)\rvert^{2}.$ It remains to show that $z^{-\frac{km}{2}}\omega(q^{-\frac{k}{2}}z)$ is nonnegative on $S^{1}$. We have
$\omega(q^{1-\frac{k}{2}}z)=q^{-\frac{m}{2}}(q^{-\frac{k}{2}}z)^{-m}\omega(q^{-\frac{k}{2}}z),$
@@
$z^{\frac{(2-k)m}{2}}\omega(q^{\frac{2-k}{2}}z)=q^{\frac{m(k-1)}{2}}z^{-\frac{km}{2}}\omega(q^{-\frac{k}{2}}z),$
so that the sign of $z^{-\frac{km}{2}}\omega(q^{-\frac{k}{2}}z)$ depends only on the parity of $k$. For $k=0,1$ we get exactly the condition in the proposition statement.
@@
$T(a\rho(a))=\int_{S^{1}}f(z)\overline{f}(z^{-1})\omega(z)\frac{dz}{z}$
should be greater than zero for all nonzero Laurent polynomials $f$ such that $f(z)=f(z^{-1})$. Using the density of trigonometric polynomials in $L^{2}(S^{1})$ and the condition $\omega(z)=\omega(z^{-1})$ we get that $\omega(z)$ should be nonnegative on $S^{1}$.
Take $a=wF(v)-w^{-1}F(v^{-1})\in\mathcal{A}$, where $F$ is a Laurent polynomial. Using (2.9) we get
$[w^{0}]a\rho(a)=q^{-\frac{m}{4}}\big{(}F(qv)v^{-\frac{m}{2}}\overline{F}(v^{-1})+F(qv^{-1})v^{\frac{m}{2}}\overline{F}(v)\big{)},$
@@
$T(a\rho(a))=q^{-\frac{m}{4}}\int_{\varepsilon S^{1}}\big{(}F(qz)z^{-\frac{m}{2}}\overline{F}(z^{-1})+F(qz^{-1})z^{\frac{m}{2}}\overline{F}(z)\big{)}\omega(z)\frac{dz}{z}.$
Using $\omega(z)=\omega(z^{-1})$ we get
$\int_{\varepsilon S^{1}}F(qz^{-1}z^{\frac{m}{2}}\overline{F}(z)\omega(z)\frac{dz}{z}=\int_{\varepsilon^{-1}S^{1}}F(qz)z^{-\frac{m}{2}}\overline{F}(z^{-1})\omega(z)\frac{dz}{z}.$
Shifting contours we get
$T(a\rho(a))=2q^{-\frac{m}{4}}\int_{\sqrt{q^{-1}}S^{1}}F(qz)\overline{F}(z^{-1})z^{-\frac{m}{2}}\omega(z)\frac{dz}{z}=2\int_{S^{1}}F(\sqrt{q}t)\overline{F}(\sqrt{q}t^{-1})t^{-\frac{m}{2}}\omega(\sqrt{q^{-1}}t)\frac{dt}{t}.$
Using the fact that trigonometric polynomials are dense in $L^{2}(S^{1})$ we get that $t^{-\frac{m}{2}}\omega(\sqrt{q^{-1}}t)$ is nonnegative on $S^{1}$. The proposition follows. ∎
Theorem 2.3.4.
Positive traces on $\mathcal{A}$ are in one-to-one correspondence with function holomorphic functions $w$ on $\mathbb{C}^{\times}$ such that
  1. 1.
@@
  2. 2.
    $\omega(z)=\omega(z^{-1})$;
  3. 3.
@@
  4. 4.
    $\omega(z)$ and $z^{-\frac{m}{2}}\omega(\sqrt{q^{-1}}z)$ are nonnegative on $S^{1}$.
The convex cone of such functions $w$ has real dimension $\frac{m}{2}-1$. In particular, for $m=4$ there is a unique positive trace up to a constant.
@@
The first statement follows from Propositions 2.2.12.2.22.3.12.3.22.3.3.
@@
We turn to the second statement. Let $w$ be a functions satisfying conditions 1-4. Consider $\omega(z)$ such that $\omega(z)=\omega(e^{2\pi iz})$. Let $q=e^{2\pi i\tau}$. Conditions 1 and 4 mean that $w$ gives a positive trace on $q$-Weyl algebra. The set of the corresponding functions $\omega$ is described in Kly (22), Theorem 3.7: function $\omega$ satisfies conditions 1 and 4 if and only if there exist $\lambda>0$ and $\alpha_{1},\ldots,\alpha_{m}$ divided into pairs $\alpha_{i}=\overline{\alpha_{j}}$ such that $\omega(z)=\lambda\prod_{i=1}^{m}\vartheta(z-\alpha_{i})$. Here $\vartheta$ is the Jacobi theta function. It follows from Lemma 3.6 of Kly (22) that we can assume $\lvert\operatorname{Im}\alpha_{i}\rvert<\operatorname{Im}\tau$.
@@
Condition 2 is equivalent to $\omega(z)=\omega(-z)$. For $z=\tfrac{1}{2}+\tfrac{\tau}{2}-\alpha_{k}$ we get $\prod_{j=1}^{m}\vartheta(\tfrac{1}{2}+\tfrac{\tau}{2}-\alpha_{k}-\alpha_{j})=0$. So there exists $j$ such that $\alpha_{j}+\alpha_{k}\in\mathbb{Z}+\mathbb{Z}\tau$.
Assume that $\alpha_{j}\neq\alpha_{k}$ or $\overline{\alpha_{k}}$. Shifting both $\alpha_{j}$ and $\overline{\alpha_{j}}$ by an integer we get $\alpha_{j}+\alpha_{k}\in\mathbb{Z}\tau$. Since $\lvert\operatorname{Im}\alpha_{j,k}\rvert<\operatorname{Im}\tau$ we have $\alpha_{j}+\alpha_{k}\in\{-\tau,0,\tau\}$.
Assume that $\alpha_{j}+\alpha_{k}=-\tau$. Note that $\vartheta(z-\alpha_{j}-\tau)=e^{\pi i(-\tau+2z-2\alpha_{j})}\vartheta(z-\alpha_{j})$ and $\vartheta(z-\overline{\alpha_{j}}+\tau)=e^{-\pi i(\tau+2z-2\overline{\alpha_{j}})}\vartheta(z-\overline{\alpha_{j}})$, so that
$\vartheta(z-\alpha_{j}-\tau)\vartheta(z-\overline{\alpha_{j}}+\tau)=e^{-2\pi i(\tau+\alpha_{j}-\overline{\alpha_{j}})}\vartheta(z-\alpha_{j})\vartheta(z-\overline{\alpha_{j}}).$
Hence we can change $\alpha_{j}$ and $\overline{\alpha_{j}}$ to $\alpha_{j}+\tau$ and $\overline{\alpha_{j}+\tau}=\overline{\alpha_{j}}-\tau$. This will multiply $\lambda$ by a positive number.
After this change we have $\alpha_{j}+\alpha_{k}=0$. Similarly, if $\alpha_{j}+\alpha_{j}=\tau$ we can shift $\alpha_{j}$ to get $\alpha_{j}+\alpha_{k}=0$.
Hence in the case when $\alpha_{j}+\alpha_{k}\in\mathbb{Z}+\mathbb{Z}\tau$ and $\alpha_{j}\neq\alpha_{k},\overline{\alpha_{k}}$ we get a quadruple $\alpha_{k},\overline{\alpha_{k}},-\alpha_{k},-\overline{\alpha_{k}}$.
If $\alpha_{j}=\alpha_{k}$ we get $\alpha_{k}\in\tfrac{1}{2}\mathbb{Z}+\frac{\tau}{2}\mathbb{Z}$. Hence $\alpha_{k}+\overline{\alpha_{k}}\in\mathbb{Z}$.
If $\alpha_{j}=\overline{\alpha_{k}}$ we get $\alpha_{k}+\overline{\alpha_{k}}\in\mathbb{Z}+\mathbb{Z}\tau$. Since $\alpha_{k}+\overline{\alpha_{k}}$ is a real number this means $\alpha_{k}+\overline{\alpha_{k}}\in\mathbb{Z}$.
@@
We get that $\alpha_{1},\cdots,\alpha_{m}$ are divided in quadruples $\alpha_{j},-\alpha_{j},\overline{\alpha_{j}},-\overline{\alpha_{j}}$ and pairs $\alpha_{k},\overline{\alpha_{k}}$ with $\operatorname{Re}\alpha_{k}\in\tfrac{1}{2}\mathbb{Z}$. Conversely, for such sequence $\alpha_{j}$ the function $\omega(z)=\lambda\prod_{i=1}^{m}\vartheta(z-\alpha_{i})$ satisfies $\omega(z)=\omega(-z)$.
The condition 3 just means that two of the pairs are ±$\pm\frac{\tau}{2}$ and $\frac{1}{2}\pm\frac{\tau}{2}$.
@@
After that, each quadruple has one complex degree of freedom and each tuple has one real degree of freedom, giving real dimension $\frac{m-4}{2}=\frac{m}{2}-2$. A choice of $\lambda$ is an additional degree of freedom. So we get a cone of real dimension $\frac{m}{2}-1$, as claimed.

### v2

@@ 9 Positive traces
$I(f_{k})=\int_{\varepsilon S^{1}}f_{k}\bigl(q^{k}z\bigr)z^{-\frac{km}{2}}\overline{f_{k}}\bigl(z^{-1}\bigr)\omega(z)\frac{{\rm d}z}{z}$
is positive for any $k$ such that $f_{k}$ is nonzero. Let $f=f_{k}$, this is a function with the poles of order at most one at $\pm q^{k}$, $k\in\mathbb{Z}$. Then the only possible poles of $f\bigl(q^{k}z\bigr)\overline{f}\bigl(z^{-1}\bigr)$ are poles of order at most two at $\pm q^{k}$, $k\in\mathbb{Z}$. Since $\omega(1)=\omega(-1)=0$ and $\omega(z)=\omega\bigl(z^{-1}\bigr)$, the function $\omega$ has a double root at $1$ and 1-1. Using the quasi-periodicity condition, we get that $w$ has a double root at $\pm q^{k}$, $k\in\mathbb{Z}$. It follows that $f\bigl(q^{k}z\bigr)\overline{f}\bigl(z^{-1}\bigr)\omega(z)$ is a holomorphic function on $\mathbb{C}^{*}$. Hence, we can write
$I(f)$ $=\int_{\varepsilon S^{1}}f\bigl(q^{k}z\bigr)z^{-\frac{km}{2}}\overline{f}\bigl(z^{-1}\bigr)\omega(z)\frac{{\rm d}z}{z}=\int_{q^{-\frac{k}{2}}S^{1}}f\bigl(q^{k}z\bigr)z^{-\frac{km}{2}}\overline{f}\bigl(z^{-1}\bigr)\omega(z)\frac{{\rm d}z}{z}$
$=q^{\frac{k^{2}m}{4}}\int_{S^{1}}f\bigl(q^{\frac{k}{2}}t\bigr)\overline{f}\bigl(q^{\frac{k}{2}}t^{-1}\bigr)t^{-\frac{km}{2}}\omega\bigl(q^{-\frac{k}{2}}t\bigr)\frac{{\rm d}t}{t}.$
For $z\in S^{1}$, we have $f\bigl(q^{\frac{k}{2}}t\bigr)\overline{f}\bigl(q^{\frac{k}{2}}t^{-1}\bigr)=\big\lvert f\bigl(q^{\frac{k}{2}}t\bigr)\big\rvert^{2}$. It remains to show that $z^{-\frac{km}{2}}\omega\bigl(q^{-\frac{k}{2}}z\bigr)$ is nonnegative on $S^{1}$. We have
$\omega\bigl(q^{1-\frac{k}{2}}zz\bigr)=q^{-\frac{m}{2}}\bigl(q^{-\frac{k}{2}}z\bigr)^{-m}\omega\bigl(q^{-\frac{k}{2}}z\bigr),$
@@
$z^{\frac{(2-k)m}{2}}\omega\bigl(q^{\frac{2-k}{2}}z\bigr)=q^{\frac{m(k-1)}{2}}z^{-\frac{km}{2}}\omega\bigl(q^{-\frac{k}{2}}z\bigr),$
so that the sign of $z^{-\frac{km}{2}}\omega\bigl(q^{-\frac{k}{2}}z\bigr)$ depends only on the parity of $k$. For $k=0,1$, we get exactly the condition in the proposition statement.
@@
$T(a\rho(a))=\int_{S^{1}}f(z)\overline{f}\bigl(z^{-1}\bigr)\omega(z)\frac{{\rm d}z}{z}$
should be greater than zero for all nonzero Laurent polynomials $f$ such that $f(z)=f\bigl(z^{-1}\bigr)$. Using the density of trigonometric polynomials in $L^{2}\bigl(S^{1}\bigr)$ and the condition $\omega(z)=\omega\bigl(z^{-1}\bigr)$, we get that $\omega(z)$ should be nonnegative on $S^{1}$.
Take $a=wF(v)-w^{-1}F\bigl(v^{-1}\bigr)\in\mathcal{A}$, where $F$ is a Laurent polynomial. Using (9.3), we get
$\big[w^{0}\big]a\rho(a)=q^{-\frac{m}{4}}\bigl(F(qv)v^{-\frac{m}{2}}\overline{F}\bigl(v^{-1}\bigr)+F\bigl(qv^{-1}\bigr)v^{\frac{m}{2}}\overline{F}(v)\bigr),$
@@
$T(a\rho(a))=q^{-\frac{m}{4}}\int_{\varepsilon S^{1}}\bigl(F(qz)z^{-\frac{m}{2}}\overline{F}\bigl(z^{-1}\bigr)+F\bigl(qz^{-1}\bigr)z^{\frac{m}{2}}\overline{F}(z)\bigr)\omega(z)\frac{{\rm d}z}{z}.$
Using $\omega(z)=\omega\bigl(z^{-1}\bigr)$, we get
$\int_{\varepsilon S^{1}}F\bigl(qz^{-1}\bigr)z^{\frac{m}{2}}\overline{F}(z)\omega(z)\frac{{\rm d}z}{z}=\int_{\varepsilon^{-1}S^{1}}F(qz)z^{-\frac{m}{2}}\overline{F}\bigl(z^{-1}\bigr)\omega(z)\frac{{\rm d}z}{z}.$
Shifting contours, we get
$T(a\rho(a))={}$ $2q^{-\frac{m}{4}}\int_{\sqrt{q^{-1}}S^{1}}F(qz)\overline{F}\bigl(z^{-1}\bigr)z^{-\frac{m}{2}}\omega(z)\frac{{\rm d}z}{z}$
$={}$ $2\int_{S^{1}}F(\sqrt{q}t)\overline{F}\bigl(\sqrt{q}t^{-1}\bigr)t^{-\frac{m}{2}}\omega\bigl(\sqrt{q^{-1}}t\bigr)\frac{{\rm d}t}{t}.$
Using the fact that trigonometric polynomials are dense in $L^{2}\bigl(S^{1}\bigr)$, we get that $t^{-\frac{m}{2}}\omega\bigl(\sqrt{q^{-1}}t\bigr)$ is nonnegative on $S^{1}$. The proposition follows. ∎
Theorem 9.4.
Positive traces on $\mathcal{A}$ are in one-to-one correspondence with holomorphic functions $\omega$ on $\mathbb{C}^{\times}$ such that
  1. $(1)$
@@
  2. $(2)$
    $\omega(z)=\omega\bigl(z^{-1}\bigr)$;
  3. $(3)$
@@
  4. $(4)$
    $\omega(z)$ and $z^{-\frac{m}{2}}\omega\bigl(\sqrt{q^{-1}}z\bigr)$ are nonnegative on $S^{1}$.
The convex cone of such functions $\omega$ has real dimension $\frac{m}{2}-1$. In particular, for $m=4$ there is a unique positive trace up to a constant.
@@ 4. $(4)$
Let $T$ be a positive trace. Using Proposition 9.2 we get that $T$ vanishes on $\mathcal{A}_{1}$. Then, by Propositions 8.1 and 8.2,
$T\left(\sum_{i=-N}^{N}w^{i}f_{i}\right)=\int f_{0}(z)\omega(z)\frac{{\rm d}z}{z}$
for a function $\omega$ satisfying the first two conditions. Using Proposition 9.1, we get $\omega(1)={\omega(-1)=0}$. For such $T$ and $\omega$, Proposition 9.3 says that the fourth condition on $\omega$ is equivalent to $T$ being positive.
@@
We turn to the second statement. Let $w$ be a functions satisfying conditions (1)–(4). Consider $w(z)$ such that $w(z)=\omega\bigl({\rm e}^{2\pi{\rm i}z}\bigr)$. Let $q={\rm e}^{2\pi{\rm i}\tau}$. Conditions (1) and (4) mean that $w$ gives a positive trace on $q$-Weyl algebra. The set of the corresponding functions $w$ is described in [17, Theorem 3.7]. Namely, the function $w$ satisfies conditions (1) and (4) if and only if there exist $\lambda>0$ and $\alpha_{1},\ldots,\alpha_{m}$ divided into pairs $\alpha_{i}=\overline{\alpha_{j}}$ such that
$w(z)=\lambda\prod_{i=1}^{m}\vartheta(z-\alpha_{i}).$ (9.4)
Here $\vartheta$ is the Jacobi theta function. It follows from [17, Lemma 3.6] that we can assume $\lvert\operatorname{Im}\alpha_{i}\rvert<\operatorname{Im}\tau$. Note that $\tau$ in this article is the same as $\tau$ in [17], but our $q$ corresponds to $q^{2}$ in [17].
Condition (2) is equivalent to w(z)=w(z)w(z)=w(-z). Evaluating at $z=\tfrac{1}{2}+\tfrac{\tau}{2}-\alpha_{k}$, we get
$\prod_{j=1}^{m}\vartheta\bigl(\tfrac{1}{2}+\tfrac{\tau}{2}-\alpha_{k}-\alpha_{j}\bigr)=0.$
@@
Hence, from any $k$ from $1$ to $m$, there exists $j$ such that $\alpha_{j}+\alpha_{k}\in\mathbb{Z}+\mathbb{Z}\tau$.
@@
We claim that we can change $\lambda,\alpha_{1},\ldots,\alpha_{m}$ so that (9.4) is still satisfied and $\alpha_{1},\ldots,\alpha_{m}$ are divided into quadruples $\alpha_{j}$, $\alpha_{j}$, $\overline{\alpha_{j}}$, $\overline{\alpha_{j}}$ and pairs $\alpha_{k}$, $\overline{\alpha_{k}}$ with $\operatorname{Re}\alpha_{k}\in\tfrac{1}{2}\mathbb{Z}$.
Let $j$, $k$ satisfy $\alpha_{j}+\alpha_{k}\in\mathbb{Z}+\mathbb{Z}\tau$. There are three cases:
  1. $(1)$
    Assume that $\alpha_{j}\neq\alpha_{k}$ or $\overline{\alpha_{k}}$. Shifting $\alpha_{j}$ by an integer we get $\alpha_{j}+\alpha_{k}\in\mathbb{Z}\tau$. Since $\lvert\operatorname{Im}\alpha_{j,k}\rvert<\operatorname{Im}\tau$ we have $\alpha_{j}+\alpha_{k}\in\{-\tau,0,\tau\}$.
    Assume that $\alpha_{j}+\alpha_{k}=-\tau$. Note that $\vartheta(z-\alpha_{j}-\tau)={\rm e}^{\pi{\rm i}(-\tau+2z-2\alpha_{j})}\vartheta(z-\alpha_{j})$ and $\vartheta(z-\overline{\alpha_{j}}+\tau)={\rm e}^{-\pi{\rm i}(\tau+2z-2\overline{\alpha_{j}})}\vartheta(z-\overline{\alpha_{j}})$, so that
    $\vartheta(z-\alpha_{j}-\tau)\vartheta(z-\overline{\alpha_{j}}+\tau)={\rm e}^{-2\pi{\rm i}(\tau+\alpha_{j}-\overline{\alpha_{j}})}\vartheta(z-\alpha_{j})\vartheta(z-\overline{\alpha_{j}}).$
    Hence, we can change $\alpha_{j}$ and $\overline{\alpha_{j}}$ to $\alpha_{j}+\tau$ and $\overline{\alpha_{j}+\tau}=\overline{\alpha_{j}}-\tau$. This will multiply $\lambda$ by a positive number. After this change, we have $\alpha_{j}+\alpha_{k}=0$. Similarly, if $\alpha_{j}+\alpha_{k}=\tau$, we can shift $\alpha_{j}$ to get $\alpha_{j}+\alpha_{k}=0$. Hence, in the case when $\alpha_{j}+\alpha_{k}\in\mathbb{Z}+\mathbb{Z}\tau$ and $\alpha_{j}\neq\alpha_{k},\overline{\alpha_{k}}$, we get a quadruple $\alpha_{k}$, $\overline{\alpha_{k}}$, $\alpha_{k}$, $\overline{\alpha_{k}}$.
  2. $(2)$
    If $\alpha_{j}=\alpha_{k}$ we get $\alpha_{k}\in\tfrac{1}{2}\mathbb{Z}+\frac{\tau}{2}\mathbb{Z}$. Hence, $\alpha_{k}+\overline{\alpha_{k}}\in\mathbb{Z}$.
  3. $(3)$
    If $\alpha_{j}=\overline{\alpha_{k}}$, we get $\alpha_{k}+\overline{\alpha_{k}}\in\mathbb{Z}+\mathbb{Z}\tau$. Since $\alpha_{k}+\overline{\alpha_{k}}$ is a real number, this means $\alpha_{k}+\overline{\alpha_{k}}\in\mathbb{Z}$.
We get that $\alpha_{1},\dots,\alpha_{m}$ are divided in quadruples $\alpha_{j}$, $\alpha_{j}$, $\overline{\alpha_{j}}$, $\overline{\alpha_{j}}$ and pairs $\alpha_{k}$, $\overline{\alpha_{k}}$ with $\operatorname{Re}\alpha_{k}\in\tfrac{1}{2}\mathbb{Z}$. Conversely, for such sequence $\alpha_{j}$ the function $\omega(z)=\lambda\prod_{i=1}^{m}\vartheta(z-\alpha_{i})$ satisfies w(z)=w(z)w(z)=w(-z).
The condition (3) just means that two of the pairs are $\pm\frac{\tau}{2}$ and $\frac{1}{2}\pm\frac{\tau}{2}$.
@@ 3. $(3)$
After that, each quadruple has one complex degree of freedom and each tuple has one real degree of freedom, giving real dimension $\frac{m-4}{2}=\frac{m}{2}-2$. A choice of $\lambda$ is an additional degree of freedom. So we get a cone of real dimension $\frac{m}{2}-1$, as claimed. ∎
Acknowledgements
The first part of the article was written by J.V. under the mentorship of D.K. at Research Science Institute, hosted by the Massachusetts Institute of Technology. We would like to thank Pavel Etingof for suggesting this project and for comments on the earlier version of the paper. We would also like to thank RSI Head Mentor, Dr. Tanya Khovanova for comments on the earlier versions of the paper. J.V. also thanks Peter Gaydarov for helpful suggestions in preparing this paper, and Allen Lin and Professor John Rickert for their thoughtful comments on an earlier version of this paper. D.K. would like to thank Davide Gaiotto for suggesting the problem discussed in the second part and to Artem Kalmykov for providing the SAGE code for computing multiplication in the algebra $\mathcal{A}$. D.K. started working on the article when he was a graduate student and then a summer research specialist at MIT. We thank the anonymous referees for thoughtful comments on this paper.